The 200- turn primary coil of a step‑down transformer is connected to an AC line that is 120.0 V (rms). The secondary coil is to supply 10.0 A (rms) at 11.4 V (rms). Assuming no power loss in the transformer, what is the number of turns ?s in the secondary coil?

?s= turns

What is the rms current ?p,rms in the primary coil?
Answer #1

Given that

N1 = 200

V1 = 120 V

I2 = 10 A

V2 = 11.4 V

Now we know that

N1/N2 = V1/V2 = I2/I1

a)

200/N2 = 120/11.4

N2 = 200*11.4/120 = 19 turns

b) 120/11.4 = 10/I1

I1 = 0.95 amp